Mr.L has 20 students in his class and 12 of them are girls, what percent are girls?

Found 2 solutions by fion6002, macston:
Answer by fion6002(19)

12/20 * 100
= 3/5 * 100
= 3 * 20
=60%

Answer by macston(5194)

(12/20)(100%)=60% ANSWER 60% of the class is girls.